OpenStudy (anonymous):

Graph the given parabola. y=x^2-6x

OpenStudy (anonymous):

please show me how to find the points

OpenStudy (anonymous):

y+9 = (x-3)^2

OpenStudy (anonymous):

upward parabola...

OpenStudy (anonymous):

with vertex 3, -9

OpenStudy (anonymous):

do u get it?

OpenStudy (anonymous):

y+9 = (x-3)^2 how did you get that?

OpenStudy (anonymous):

u add 9 on both sides

OpenStudy (anonymous):

why 9?

OpenStudy (anonymous):

on the RHS u get x^2 - 6x +9 which is the square of (x-3)

OpenStudy (anonymous):

ohhh i see

OpenStudy (anonymous):

so u get the eqn....

OpenStudy (anonymous):

yea thx!

OpenStudy (anonymous):

so tell me where the focus is...

OpenStudy (anonymous):

(0,0) (6,0)

OpenStudy (anonymous):

focus

OpenStudy (anonymous):

focus?

OpenStudy (anonymous):

whts the coordinates of the focus?

OpenStudy (anonymous):

(3, -35/4)

OpenStudy (anonymous):

hmmm